Consider the following code mapping for k = 2 bits of information with n = 5. How many errors at maximum can this code correct?

What are the generator and parith check matrices of the code? (Enough to provide one valid parity check matrix).

Information bits

Codeword

00

00000

01

00111

10

11100

11

11011
Answer #1

Ans:-

Part A It can correct at most 1 error because minimum hamming distance is 3 and maximum errors corrected = (HD-1)/2 = (3-1)/2

NOTE:- If u have any doubts please comment below and I am happily help to u brother please thumsup bro and don't give any negative reviews please ????